This is page 1

Printer: Opaque this

Put Your Title Here

author 1 author 2 etc.

ABSTRACT This is a small sample document which should help you in formatting your paper for publication in the Case Studies in Bayesian Statistics volume. If you have any questions, you can contact me via email at heidi@stat.cmu.edu or by phone (412) 268-2718. I hope you will find this helpful.

1 Introduction

2 Using this template

Using this template and the svcon.sty file should make setting up your paper in Springer-Verlag style very easy. It is already formatted for margins and page size, etc. You should start by filling in the chapter, shortchapname and chapterauthors commands above. Then insert your text, putting section and subsection titles as shown.

2.1 Equations

Here is an equation, showing how to use the “label” and “ref” function:

2 + 2 = 4  (2.1)

Now we have another equation.

5 − 5 = 0,  (2.2)

which we will reference a bit later.

Some more text, and another equation.

7 + 7 = 14 ,  (2.3)

after which we might refer to the second equation, as in, “as can be seen from equation (??).” (If (??) appears instead of the equation number, just re-L^{A}TEXthe file.)

3 More Hints

3.1 Reference style

Here are two examples of the reference style we are using. Experimental evidence supporting the first hypothesis has been provided by several authors (Rudolph et al., 1967; Heyman et al., 1977; von Ritter et al., 1988). For a precise description of these models we refer to Almeida and Gidas (1993).

3.2 Figures

Figures can be included as .ps files. Psfig is an efficient way of handling postscript files. There are many sources of info about Psfig on the World Wide Web. If you do not have the psfig.sty file available to you, please let me know and I will email it to you.

4 Final Steps

When your paper is in final form, please email me the L^{A}TEX file (don’t forget any included .ps files) and send a hard copy. My mailing address is Heidi Sestrich, Dept. of Statistics, Baker Hall 232A, Carnegie Mellon University, Pittsburgh, PA 15213, email is heidi@stat.cmu.edu.

Remember that we are preparing a camera ready copy for Springer-Verlag. The published paper will look *exactly* the same as the paper you send me.

Acknowledgments

This paper was supported by a grant from the National Science Foundation. The authors would like to thank Albert Einstein for riding a bicycle.

References

Almeida, M. and Gidas, B. (1993). A Variational Method for Estimating the Parameters of MRF from Complete or Incomplete Data, Annals Appl. Prob. 3, 103-136.

Heyman, M.A., Payne B.D., Hoffman, J.I.E. and Rudolph, A.M. (1977). Blood flow measurements with radionuclide-labled particles. Progress in Cardiovascular Diseases 10, 55-79.